# Imprints of an early matter-dominated era arising from dark matter dilution mechanism on cosmic string dynamics and gravitational wave signatures

### v2

@@ IV Summary
In this work, we explore the influence of an EMD era on the dynamics of a preexisting CS network and the resulting SGWB. We first review the GWs generated by CS loops, highlighting the explicit dependence of the SGWB spectrum on the CS loop number density $n_{\mathrm{CS}}$ and the scale factor $a(t)$. We then analyze the evolution of the CS network based on the VOS model and discuss the scaling behavior. By assuming appropriate forms for the loop production functions, we derive expressions for the CS loop number densities in RD and MD eras, which are consistent with numerical simulations in the scaling regime. Notably, these expressions are expected to remain valid even in cases where the scaling behavior is violated.
@@
Furthermore, we show that the evolution of the normalized correlation length $\xi$ and the RMS velocity $v$ of the CS network is modified by the presence of the EMD era, exhibiting a nonscaling behavior. Subsequently, we calculate the number density of CS loops formed during the EMD era, as well as the number densities of loops generated in earlier eras and surviving into later eras. Using these results, the influence of the EMD era on the SGWB spectrum arising from CS loops is demonstrated.

### v3

@@ IV Summary
In this study, we explore the influence of an EMD era on the dynamics of a preexisting CS network and the resulting SGWB. We first review the GWs generated by CS loops, highlighting the explicit dependence of the SGWB spectrum on the CS loop number density $n_{\mathrm{CS}}$ and the scale factor $a(t)$. We then analyze the evolution of the CS network based on the VOS model and discuss the scaling behavior. By assuming appropriate forms for the loop production functions, we derive expressions for the CS loop number densities in RD and MD eras, which are consistent with numerical simulations in the scaling regime. Notably, these expressions are expected to remain valid even in cases where the scaling behavior is violated.
@@
Furthermore, we show that the evolution of the normalized correlation length $\xi$ and the RMS velocity $v$ of the CS network is modified by the presence of the EMD era, exhibiting nonscaling behavior. Subsequently, we calculate the number density of CS loops formed during the EMD era, as well as the number densities of loops generated in earlier eras and surviving into later eras. Using these results, the influence of the EMD era on the SGWB spectrum arising from CS loops is demonstrated.